# Analytic and Variational Stability of Deep Learning Systems

### v1

@@ 5 Residual and Feedforward Stability Laws
We now specialise the analytic stability theory to standard architectures. We consider feedforward networks with spectral norm control and residual networks with an explicit step size. The purpose is not to model all implementation details, but to extract simple analytic laws that recover and sharpen common stability heuristics.
@@ 5.1 Feedforward networks with spectral control
Consider an $L$-layer feedforward network of the form
@@
where $W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}}$ and $b_{k}\in\mathbb{R}^{d_{k+1}}$. Assume that $\sigma$ is applied elementwise, is $1$-Lipschitz, and is differentiable almost everywhere. ReLU, leaky ReLU with slope in $[0,1]$, and hard $\tanh$ satisfy these properties.
At differentiable points the Jacobian has the form
@@
where $D_{k}(x)$ is diagonal with entries in $[0,1]$. Consequently,
$\|Jf_{k}(x)\|_{2}\leq\|W_{k}\|_{2}\quad\text{for all differentiable points }x.$
Let $F_{\theta}:=f_{L-1}\circ\cdots\circ f_{0}$, and let $\mathsf{S}^{x}_{L}$ denote the forward sensitivity operator in the sense of Definition 3.1. By the chain rule,
@@
Consider the feedforward architecture above and suppose there exists $\rho\in(0,1)$ such that
@@
In particular, the forward component of the Learning Stability Profile is uniformly bounded, and the hypotheses of Theorem 4.1 hold with a quadratic energy depending only on $X_{L}$ and $\theta$.
@@
The Jacobian bound implies that the forward Lipschitz constant of each layer does not exceed $\rho$. Applying the chain rule yields
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|Jf_{k}(X_{k})\|_{2}\leq\rho^{L}.$
This gives the uniform bound on the forward sensitivity and implies that the forward stability signatures decay exponentially with depth. The exponent bound follows directly from the definition of $\alpha_{x}$, and existence of a compatible energy follows from Theorem 4.1. ∎
Theorem 5.1 formalises the standard intuition that spectral norm control ensures stability of deep feedforward networks [Bartlett2017Spectrally, Yoshida2017]. Within the analytic stability framework, the forward exponent $\alpha_{x}$ acts as a depth-normalised logarithmic Lipschitz constant, and the condition $\rho<1$ enforces strict contraction along the representation direction.
@@ 5.2 Residual networks and CFL type stability
Residual architectures introduce an explicit step size that links depthwise propagation with time discretisations of continuous dynamical systems [He2016ResNet, HaberRuthotto2017]. Within the analytic stability framework this connection becomes quantitative through the forward stability exponent.
@@
Assume that $g_{k}$ is globally Lipschitz in $x$ with constant $M_{g}$ and uniformly dissipative with rate $m>0$, meaning
$\frac{G_{k}(x;\theta_{k})+G_{k}(x;\theta_{k})^{\top}}{2}\preceq-mI,\qquad\|G_{k}(x;\theta_{k})\|_{2}\leq M_{g}$ (5.2)
for all admissible $(x,\theta_{k})$ and all $k$. Such conditions are standard in the analysis of dissipative ODEs and their explicit time discretisations [Strikwerda2004, HaberRuthotto2017].
At a fixed layer the Jacobian with respect to $X_{k}$ is
@@
Let $v$ be a unit vector. Then
$\|(I+hG_{k})v\|_{2}^{2}=\|v\|_{2}^{2}+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|_{2}^{2}.$
The dissipativity assumption yields $v^{\top}G_{k}v\leq-m\|v\|_{2}^{2}$, while the Lipschitz bound gives $\|G_{k}v\|_{2}^{2}\leq M_{g}^{2}\|v\|_{2}^{2}$. Substitution gives
$\|(I+hG_{k})v\|_{2}^{2}\leq(1-2hm+h^{2}M_{g}^{2})\|v\|_{2}^{2}.$
Taking the supremum over unit vectors proves the claim. ∎
@@
For sufficiently small $h$, this factor is strictly less than one and governs depthwise propagation of perturbations. The next result translates this local bound into a global constraint on the forward analytic exponent.
@@
Consider the residual architecture (5.1) under assumption (5.2). Suppose the step size satisfies
$0<h<h_{\max},\qquad h_{\max}:=\frac{2m}{M_{g}^{2}}.$ (5.3)
@@
In particular, the forward component of the Learning Stability Profile is uniformly bounded in depth, and the hypotheses of Theorem 4.1 hold with a quadratic energy in $X_{k}$.
@@
By Lemma 5.2, each residual layer has spectral norm at most $c_{x}(h)$. The composition of $L$ layers therefore has forward Jacobian norm bounded by $c_{x}(h)^{L}$. If $h<h_{\max}$ then $1-2hm+h^{2}M_{g}^{2}<1$ and hence $c_{x}(h)<1$. The bound on $\alpha_{x}$ follows directly from Definition 3.1. Existence of a compatible learning energy then follows from Theorem 4.1. ∎
Condition (5.3) is a discrete stability restriction. Its structure is identical to a Courant–Friedrichs–Lewy type condition for explicit time stepping schemes applied to dissipative systems [Strikwerda2004]. The dimensionless quantity $hM_{g}$ measures the scale of each residual perturbation relative to intrinsic dissipation. If $h$ is too large, the contraction factor $c_{x}(h)$ exceeds one and the forward exponent becomes positive. The analytic stability framework therefore formalises the heuristic that residual blocks must be either sufficiently small in step size or strongly dissipative to ensure stable propagation [He2016ResNet, HaberRuthotto2017].
In many practical settings the parameters $\theta_{k}$ are updated by an inner gradient loop. Under the smoothness and convexity assumptions introduced in Section 6, such updates induce a parametric Lipschitz factor strictly smaller than one. Combining Theorem 5.3 with the stochastic gradient analysis yields negative forward and parametric exponents $(\alpha_{x},\alpha_{\theta})$, with $\alpha_{u}$ controlled by the noise level and learning rate. Residual networks satisfying the spectral and step-size conditions therefore lie in the analytically stable regime of Theorem 4.1.
@@
In this regime the forward analytic exponent satisfies
$\alpha_{x}\leq\frac{1}{2}\log\bigl(1-2hm+h^{2}M_{g}^{2}\bigr)<0,$
so the representation dynamics are strictly contractive across depth.
@@ 6 Temporal stability of stochastic gradient methods
We consider stochastic gradient updates of the form
$\theta^{t+1}=\theta^{t}-\eta_{t}\,G(\theta^{t},u^{t}),\qquad t\geq 0,$ (6.1)
where $(\eta_{t})_{t\geq 0}$ is a step-size sequence and $(u^{t})_{t\geq 0}$ is an i.i.d. source of stochasticity.
@@
The objective $\mathcal{L}$ is $\mu$–strongly convex and LL–smooth. The stochastic gradient oracle satisfies
$\mathbb{E}[G(\theta,u)\mid\theta]=\nabla\mathcal{L}(\theta),$
and there exist constants $\sigma_{0},\sigma_{1}\geq 0$ such that
$\mathbb{E}\!\left[\|G(\theta,u)-\nabla\mathcal{L}(\theta)\|^{2}\mid\theta\right]\leq\sigma_{0}^{2}+\sigma_{1}^{2}\|\theta-\theta^{\star}\|^{2}.$
Moreover, there exist constants $L_{G},L_{u}^{G}>0$ such that for all $\theta,\bar{\theta}\in\Theta$ and $u,\bar{u}\in\mathcal{U}$,
$\|G(\theta,u)-G(\bar{\theta},u)\|\leq L_{G}\|\theta-\bar{\theta}\|,\qquad\|G(\theta,u)-G(\theta,\bar{u})\|\leq L_{u}^{G}\|u-\bar{u}\|.$ (6.2)
@@ 6.1 Mean square Lyapunov recursion
We derive a mean square Lyapunov recursion for the parameter error. Define $e^{t}:=\theta^{t}-\theta^{\star}$ and decompose the stochastic gradient as
$G(\theta^{t},u^{t})=\nabla\mathcal{L}(\theta^{t})+\zeta^{t},\qquad\zeta^{t}:=G(\theta^{t},u^{t})-\nabla\mathcal{L}(\theta^{t}).$
By Assumption 3, $\mathbb{E}[\zeta^{t}\mid\theta^{t}]=0$ and
$\mathbb{E}\bigl[\|\zeta^{t}\|^{2}\mid\theta^{t}\bigr]\leq\sigma_{0}^{2}+\sigma_{1}^{2}\|e^{t}\|^{2}.$
The update (6.1) becomes
$e^{t+1}=e^{t}-\eta_{t}\nabla\mathcal{L}(\theta^{t})-\eta_{t}\zeta^{t}.$
Conditioning on $\theta^{t}$ and expanding yields
$\mathbb{E}\bigl[\|e^{t+1}\|^{2}\mid\theta^{t}\bigr]=\|e^{t}-\eta_{t}\nabla\mathcal{L}(\theta^{t})\|^{2}+\eta_{t}^{2}\mathbb{E}\bigl[\|\zeta^{t}\|^{2}\mid\theta^{t}\bigr].$
Strong convexity and smoothness of $\mathcal{L}$ imply
$\langle\nabla\mathcal{L}(\theta^{t}),e^{t}\rangle\geq\mu\|e^{t}\|^{2},\qquad\|\nabla\mathcal{L}(\theta^{t})\|^{2}\leq L^{2}\|e^{t}\|^{2}$
for all $\theta^{t}$ [Nesterov2004, Chapter 2]. Consequently,
$\|e^{t}-\eta_{t}\nabla\mathcal{L}(\theta^{t})\|^{2}\leq\bigl(1-2\eta_{t}\mu+\eta_{t}^{2}L^{2}\bigr)\|e^{t}\|^{2}.$
Combining the above estimates yields the Lyapunov recursion
$\mathbb{E}\bigl[\|e^{t+1}\|^{2}\mid\theta^{t}\bigr]\leq q_{t}\|e^{t}\|^{2}+\eta_{t}^{2}\sigma_{0}^{2},$ (6.3)
where
$q_{t}:=1-2\eta_{t}\mu+\eta_{t}^{2}(L^{2}+\sigma_{1}^{2}).$
This inequality is the fundamental mean square energy estimate for the parameter energy
$\mathcal{E}(\theta):=\|\theta-\theta^{\star}\|^{2}.$
where the cross term vanishes since $\mathbb{E}[\zeta^{t}\mid\theta^{t}]=0$ by Assumption 3. The preceding computation can be summarized as the following lemma.
@@
Let Assumption 3 hold and let $\mathcal{E}(\theta)=\|\theta-\theta^{\star}\|^{2}$. Then the SGD update (6.1) satisfies
$\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\mid\theta^{t}\bigr]\leq\mathcal{E}(\theta^{t})-\eta_{t}\bigl(2\mu-\eta_{t}(L^{2}+\sigma_{1}^{2})\bigr)\mathcal{E}(\theta^{t})+\eta_{t}^{2}\sigma_{0}^{2}$
for all $t\geq 0$.
@@ 6.2 Temporal analytic stability for constant and decreasing step sizes
    The quadratic energy $\mathcal{E}(\theta)=\|\theta-\theta^{\star}\|^{2}$ obeys
@@
    for all $t$. In particular, the mean square error converges to a stationary noise floor of order $\eta$.
@@
Part (1) follows directly from (6.3) and the tower property of conditional expectation. Under (6.4), a standard linear recursion argument gives
$\mathbb{E}\bigl[\|e^{t}\|^{2}\bigr]\leq q(\eta)^{t}\|e^{0}\|^{2}+\frac{\eta^{2}\sigma_{0}^{2}}{1-q(\eta)}.$
For part (2), the Jacobian of the mean update map $\theta\mapsto\theta-\eta\nabla\mathcal{L}(\theta)$ has spectral radius at most $\sqrt{q(\eta)}$. This contraction rate governs the linearised dynamics, yielding $\alpha_{\theta}\leq\tfrac{1}{2}\log q(\eta)$ in the sense of Definition 3.1.
For part (3), differentiating the update map with respect to $u$ and using (6.2) gives an inhomogeneous linear system for perturbations driven by $\delta u^{t}$. Standard input–output bounds for linear time-varying systems [Khalil2002Nonlinear, Chapter 4] imply a uniform gain from $(\delta u^{0},\dots,\delta u^{t})$ to $\delta\theta^{t}$, yielding the stated bound and $\alpha_{u}\leq 0$. ∎
With constant step size, stability is not asymptotic in the deterministic sense, since the iterates fluctuate around $\theta^{\star}$ at a scale determined by $\eta$ and the noise variance. The parametric analytic exponent $\alpha_{\theta}$ nevertheless captures the exponential forgetting of initial conditions in mean square, while $\alpha_{u}$ quantifies the uniform sensitivity to perturbations in the stochastic update mechanism.
We now turn to the decreasing step size regime. In that case the iterates converge almost surely to $\theta^{\star}$ and both analytic exponents are nonpositive, in agreement with classical stochastic approximation theory [KushnerYin2003, Benaim1999].
@@
    The sequence $(\theta^{t})$ converges almost surely to the unique minimiser $\theta^{\star}$. The energy $\mathcal{E}(\theta^{t})$ converges almost surely and is nonincreasing in expectation.
@@
    $\alpha_{u}\leq 0.$
    More precisely, the temporal stability signatures $\sigma^{u}_{t}$ form a bounded sequence.
@@
The argument follows the classical ODE method for stochastic approximation [KushnerYin2003, Benaim1999]. Under Assumption 3, the mean dynamics associated with (6.1) track the gradient flow
$\dot{\theta}(s)=-\nabla\mathcal{L}(\theta(s)),$
which has a globally asymptotically stable equilibrium at $\theta^{\star}$ by strong convexity.
The step size conditions ensure that the stochastic perturbations are square summable while the mean drift is persistent. Standard almost supermartingale arguments (Robbins–Siegmund theorem; see [KushnerYin2003, Ch. 2]) imply that $\theta^{t}\to\theta^{\star}$ almost surely and that the Lyapunov energy $\mathcal{E}(\theta^{t})$ converges almost surely and decreases in expectation.
To obtain the analytic exponent bounds, note that convergence of $\theta^{t}$ together with the Jacobian bounds in (6.2) imply that the linearised parameter dynamics are asymptotically nonexpansive. Consequently, the parametric sensitivity operators remain uniformly bounded, yielding $\alpha_{\theta}\leq 0$.
Similarly, the Jacobian bound with respect to $u$ ensures that the linear response to perturbations of the update sequence remains uniformly bounded along the trajectory. Hence the temporal stability signatures $\sigma^{u}_{t}$ do not grow exponentially, and $\alpha_{u}\leq 0$. ∎
Theorems 6.2 and 6.3 place stochastic gradient methods squarely within the analytic stability framework developed in this paper. The quadratic energy $\mathcal{E}(\theta)$ acts as a Lyapunov function in expectation, while the learning rate conditions determine whether the parametric and temporal analytic exponents are strictly negative or merely nonpositive. In both regimes, the Learning Stability Profile provides a unified quantitative description of robustness to perturbations of initialisation and stochastic update mechanisms.
@@ 7 Variational Extension to Non-smooth Learning Systems
The analytic framework developed so far assumes that the layer maps $f_{k}$ and update maps $\mathcal{A}_{t}$ are continuously differentiable. Many architectures of practical interest violate this assumption. Common examples include ReLU and max pooling activations, hinge and $\ell_{1}$ losses, clipping and quantisation operators, and proximal or projected optimisation steps. This section extends the stability theory to such non-smooth learning systems.
The extension replaces classical Jacobians with Clarke generalized derivatives and smooth Lyapunov energies with variational energies that may be non differentiable. The resulting theory preserves the central equivalence established earlier: bounded stability signatures are equivalent to the existence of an energy that dissipates along learning trajectories. Thus the analytic exponents and stability laws remain valid in the non-smooth regime when derivatives are interpreted in the sense of Clarke [Clarke1990Optimization, Clarke1998Nonsmooth, RockafellarWets1998].
7.1 Locally Lipschitz learning flows and Clarke derivatives
We work in finite dimensional Euclidean spaces equipped with norms compatible with the product structure of the state space. The following assumption replaces Assumption 1.

### v2

@@ 4 Residual and Feedforward Stability Laws
We now specialise the analytic stability theory to standard architectures. The goal is not to capture implementation details, but to extract simple analytic conditions that recover common stability heuristics as direct consequences of the general framework.
@@ 4.1 Feedforward networks with spectral control
Consider an $L$-layer feedforward network
@@
where $\sigma$ is applied elementwise and is $1$-Lipschitz. ReLU and leaky ReLU with slope in $[0,1]$ satisfy this condition.
At differentiable points, the Jacobian takes the form
@@
where $D_{k}(x)$ is diagonal with entries in $[0,1]$. Hence
$\|Jf_{k}(x)\|_{2}\leq\|W_{k}\|_{2}.$
Let $F_{\theta}:=f_{L-1}\circ\cdots\circ f_{0}$. By the chain rule,
@@
Suppose there exists $\rho\in(0,1)$ such that
@@
Each layer has Lipschitz constant at most $\rho$. Applying the chain rule yields an overall Lipschitz constant bounded by $\rho^{L}$. The exponent bound follows directly from the definition of $\alpha_{x}$. Existence of a compatible learning energy follows from Theorem 3.1. ∎
Theorem 4.1 recovers the standard principle that spectral norm control enforces stability in deep feedforward networks. Within the present framework, the forward exponent $\alpha_{x}$ plays the role of a depth-normalised logarithmic Lipschitz constant, and the condition $\rho<1$ enforces strict contraction along the representation direction.
@@ 4.2 Residual networks and CFL type stability
Residual architectures introduce an explicit step size linking depthwise propagation to time discretisations of continuous dynamical systems He et al. [2016]; Haber and Ruthotto [2017]. Within the analytic stability framework, this connection is captured quantitatively by the forward stability exponent.
@@
Assume $g_{k}$ is globally Lipschitz and uniformly dissipative, i.e.
$\frac{G_{k}(x;\theta_{k})+G_{k}(x;\theta_{k})^{\top}}{2}\preceq-mI,\qquad\|G_{k}(x;\theta_{k})\|_{2}\leq M_{g},$ (4.2)
for all admissible $(x,\theta_{k})$ and all $k$. Such conditions are standard in the stability analysis of explicit discretisations of dissipative systems Strikwerda [2004].
At each layer,
@@
For any unit vector $v$,
$\|(I+hG_{k})v\|_{2}^{2}=1+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|_{2}^{2}.$
The dissipativity and Lipschitz bounds imply $v^{\top}G_{k}v\leq-m$ and $\|G_{k}v\|_{2}^{2}\leq M_{g}^{2}$, yielding the claim. ∎
@@
Consider the residual architecture (4.1) under assumption (4.2). If the step size satisfies
$0<h<h_{\max},\qquad h_{\max}:=\frac{2m}{M_{g}^{2}},$ (4.3)
@@
Lemma 4.2 gives a uniform bound $c_{x}(h)$ on each layer. Composition over depth yields a forward Jacobian norm bounded by $c_{x}(h)^{L}$. The exponent bound follows directly from the definition of the analytic exponent given in Section 2.
Condition (4.3) is a discrete stability restriction identical in structure to a Courant-Friedrichs-Lewy condition for explicit time stepping of dissipative systems Strikwerda [2004]. The analytic stability framework therefore formalises the heuristic that residual blocks must be either sufficiently small in step size or sufficiently dissipative to ensure stable depthwise propagation.
@@
In this regime,
$\alpha_{x}\leq\frac{1}{2}\log\bigl(1-2hm+h^{2}M_{g}^{2}\bigr)<0.$
@@ 5 Temporal stability of stochastic gradient methods
We consider stochastic gradient updates
$\theta^{t+1}=\theta^{t}-\eta_{t}\,G(\theta^{t},u^{t}),$ (5.1)
under standard strong convexity, smoothness, and bounded noise assumptions Nesterov [2004]; Kushner and Yin [2003].
We assume that the stochastic gradient map is uniformly Lipschitz in both parameters and update randomness. Namely, there exist constants $L_{G},L_{u}^{G}\geq 0$ such that
$\|G(\theta_{1},u)-G(\theta_{2},u)\|\leq L_{G}\|\theta_{1}-\theta_{2}\|,\qquad\|G(\theta,u_{1})-G(\theta,u_{2})\|\leq L_{u}^{G}\|u_{1}-u_{2}\|.$ (5.2)
@@
The objective $\mathcal{L}$ is $\mu$–strongly convex and LL–smooth. The stochastic gradient satisfies
$\mathbb{E}[G(\theta,u)\mid\theta]=\nabla\mathcal{L}(\theta),\qquad\mathbb{E}\!\left[\|G(\theta,u)-\nabla\mathcal{L}(\theta)\|^{2}\mid\theta\right]\leq\sigma_{0}^{2}+\sigma_{1}^{2}\|\theta-\theta^{\star}\|^{2}.$
Moreover, $G$ is Lipschitz in $\theta$ and $u$ as in (5.2).
@@ 5.1 Mean square Lyapunov recursion
Let $e^{t}:=\theta^{t}-\theta^{\star}$. Standard expansion yields the conditional recursion
$\mathbb{E}\bigl[\|e^{t+1}\|^{2}\mid\theta^{t}\bigr]\leq\bigl(1-2\eta_{t}\mu+\eta_{t}^{2}(L^{2}+\sigma_{1}^{2})\bigr)\|e^{t}\|^{2}+\eta_{t}^{2}\sigma_{0}^{2}.$ (5.3)
@@
Under Assumption 1, the SGD update (5.1) satisfies
$\mathbb{E}\bigl[\|\theta^{t+1}-\theta^{\star}\|^{2}\mid\theta^{t}\bigr]\leq\|\theta^{t}-\theta^{\star}\|^{2}-\eta_{t}\bigl(2\mu-\eta_{t}(L^{2}+\sigma_{1}^{2})\bigr)\|\theta^{t}-\theta^{\star}\|^{2}+\eta_{t}^{2}\sigma_{0}^{2}.$
@@ 5.2 Temporal analytic stability for constant and decreasing step sizes
    for all $t$.
@@
Part (1) follows directly from (5.3). The linear contraction of the mean update map implies exponential forgetting of initial conditions in mean square, yielding part (2). Part (3) follows from differentiating the update with respect to $u$ and applying standard input–output bounds for linear time-varying systems [Khalil, 2002, Chapter 4]. ∎
With constant step size, stability is not asymptotic in the deterministic sense due to persistent noise. The exponent $\alpha_{\theta}$ nevertheless captures exponential decay of initialization effects, while $\alpha_{u}$ quantifies uniform sensitivity to stochastic perturbations.
@@
    The iterates $\theta^{t}$ converge almost surely to $\theta^{\star}$, and the energy $\mathcal{E}(\theta^{t})$ converges almost surely and is nonincreasing in expectation.
@@
    $\alpha_{u}\leq 0,$
    and the temporal stability signatures remain uniformly bounded.
@@
The argument follows the classical ODE method for stochastic approximation Kushner and Yin [2003]; Benaïm [1999]. Almost sure convergence and boundedness of sensitivities imply nonexpansiveness of the linearised dynamics, yielding $\alpha_{\theta}\leq 0$ and $\alpha_{u}\leq 0$. ∎
Theorems 5.2 and 5.3 embed stochastic gradient methods within the analytic stability framework. The learning rate determines whether the parametric and temporal exponents are strictly negative or merely nonpositive, while the Learning Stability Profile provides a unified description of robustness to initialization and stochastic update perturbations.
@@ 6 Variational Extension to Non-smooth Learning Systems
Many learning architectures involve non-smooth components, including ReLU-type activations, nonsmooth losses, clipping, quantisation, and proximal or projected updates. This section extends the analytic stability framework to such systems.
The extension replaces classical Jacobians with Clarke generalized derivatives and smooth Lyapunov energies with variational energies. The central equivalence established earlier remains valid: bounded stability signatures are equivalent to the existence of a dissipative learning energy, with derivatives interpreted in the Clarke sense Clarke [1990]; Clarke et al. [1998]; Rockafellar and Wets [1998].
6.1 Locally Lipschitz learning flows
We replace differentiability assumptions by local Lipschitz continuity.